Translate the following sentence into an equation 1/5 of the sum of a number and 12 is 27
Answer by ewatrrr(10682)


Hi
Let x represent the number
Write as we Read***
1/5(x+12) = 27
x = 27*5 - 12
x = 123
CHECKING our Answer***
1/5(135) = 27